OpenStudy (loser66):

At 7am one morning detectives find a murder victim in a closed meat locker. The temperature of the victim measures 88degrees. Assume the meat locker is always kept at 40 degrees, and at the time of death the victim's temperature was 98.6 degrees. When the body is finally removed at 8am, its temperature is 86 degrees. a) when did the murder occur? b) How big an error in the time of death would result if the live body temperature was known only to be between 98.2 degrees and 101.4 degrees? Please, help

OpenStudy (loser66):

@dan815

OpenStudy (freckles):

i'm thinking we will use newton's law of cooling

OpenStudy (loser66):

Yes, it is. But I got negative value which is impossible. That is why I posted it here.

OpenStudy (freckles):

for question A?

OpenStudy (loser66):

Let 7:00am be the initial time, we set T(0) = 88. And the Newton's law gives us \(T(t) = 40 -\alpha e^{-kt}\) I put t =0 to solve for alpha After 1hour (8:00 am) , T(60) = 86 to solve for k Hence, I got \(T(t) = 40 +48e^{-0.000709t}\) But, if I calculate the time T(t) = 98.6, I got t = -281.... :(

OpenStudy (freckles):

well -281 would just mean 281 mins before the initial time 7am

OpenStudy (loser66):

No, it is not that since they remove the body at 8:00, So it is impossible its death happened after 8

OpenStudy (freckles):

what?

OpenStudy (freckles):

281 min before 7am is definitely before 7am or 8am

OpenStudy (loser66):

And if it is before 7:00, how can the degree increase from 88 to 98.6?

OpenStudy (freckles):

I'm going to do the problem... but from your answer it looks like you have 281 mins before 7am which is ...2:19 am

OpenStudy (freckles):

give me a sec to check your other work

OpenStudy (freckles):

\[T(t) \\ \text{ where } t \text{ is \in minutes } \\ T(0)=T_0=88^\circ \text{ ---this is at } 7am \\ T(60)=T_{60}=86^o \text{--- 60 minutes later which is } 8am \\ \\ T_a =40^\circ \text{ ---constant temp of the refrigerator } \\ \frac{dT}{dt}=-k(T-40^ \circ) \\ \frac{dT}{T-40^\circ}=-k dt \\ \ln|T-40^\circ|=-kt+C \\ T-40^\circ=a e^{-kt} \\ T=40^\circ+ae^{-kt} \\ T_0=40^\circ+a=88^\circ \implies a= 88^\circ-40^\circ=48^ \circ \\ \text{ so now we have } \\ T=40^\circ+48^\circ e^{-k t} \\ T_{60}=40^\circ+48^\circ e^{-k \cdot 60} =86^\circ \implies \\ e^{- k 60}=\frac{46}{48} \\ -k 60 =\ln(\frac{46}{48}) \\ k=\frac{-1}{60} \ln(\frac{46}{48}) \\ k=0.00070933\] so far all of your stuff is right up to this point

OpenStudy (freckles):

\[T=40^\circ+48^ \circ e^{-0.00070933 t}\] now we are given T at the time of death and asked to find t

OpenStudy (freckles):

\[98.6^\circ=40^\circ+48^\circ e^{-0.00070933 t}\] solving this will give us the answer in terms of the initial time being 7am so if the t is negative that means it happened before 7am

OpenStudy (freckles):

which would make sense because how can the death happen after them finding the body :p

OpenStudy (freckles):

and I also get t is approximately -281 when T is 98.6 which means 281 minutes before 7am is when the death happened

OpenStudy (loser66):

I am thinking of the case \(a =\pm e^C\)

OpenStudy (freckles):

\[281 \text{ minutes }=4 \text{ hours } \text{ and } 41 \text{ minutes } \\ 7 am -(4 \text{ hours and } 41 \text{ minutes })=3 am -(41 \text{minutes }) \\ =(2 \text{ hours} \text{ and } 60 \text{ minutes } ) -(41 \text{ minutes } ) \\ =2 \text{ hours } \text{ and } (60 \text{ minutes } -41 \text{ minutes }) \\ =2 \text{ hours } \text{ and } 19 \text{ minutes } \\ =2:19 am\]

OpenStudy (loser66):

We already worked on a = e^C, (a>0), but it still has the case a<0 since |T-40| has 2 values of the right hand side.

OpenStudy (loser66):

I meant from \(ln|T-40|= -kt+C\), we have \(|T-40|= e^{-kt +C}=e^Ce^{-kt}\)\) Hence if we let a = e^C, then a must be \(a =\pm e^C\)

OpenStudy (freckles):

which question are you working on?

OpenStudy (freckles):

e^c is positive

OpenStudy (loser66):

a) I am still doubt when the result is -281

OpenStudy (freckles):

why?

OpenStudy (freckles):

the negative just means before the initial time which was 7am

OpenStudy (freckles):

-281= 281 minutes before 7am=2:19am

OpenStudy (loser66):

because the Newton's of cooling applied to heating also, right?

OpenStudy (freckles):

well the body is cooling over time not heating ...

OpenStudy (freckles):

so before 7am the body should be hotter

OpenStudy (freckles):

which it is

OpenStudy (loser66):

oh, I got what you meant

OpenStudy (freckles):

http://www.sosmath.com/diffeq/first/application/newton/newton.html here is a very similar example

OpenStudy (freckles):

end of page

OpenStudy (freckles):

the - just means before the initial time you used

OpenStudy (freckles):

we should be able to use 8am as the initial time if you wanted but 7am would have to be at t=-60

OpenStudy (loser66):

:) I used their computer to see the answer, it gave me 2:22:36

OpenStudy (freckles):

2:19am is pretty close -281 was an approximated answered just like our value for k was too

OpenStudy (loser66):

Ooooooooooooook how about part b?

OpenStudy (freckles):

\[\text{ so } T=40^\circ+48^\circ e^{- .00070933t} \\ \text{ sounds like it wants us to solve for } t \\ \text{ given } 98.2^\circ<T<101.4^\circ\]... maybe

OpenStudy (freckles):

and then once you solve for t you can subtract -281 on all sides just to show how close to the actual time the numbers were

OpenStudy (loser66):

Yes, I think so. If it is so, we just calculate t for the limits only and then subtract to the previous result, right?

OpenStudy (loser66):

I got what you mean. Thanks for the help. :)

OpenStudy (freckles):

np and yes you solve both T=98.2 and T=101.4 separately because we know this is a decreasing function (there are no weird increases to worry about)

OpenStudy (loser66):

:)